# DiffCast: A Unified Framework via Residual Diffusion for Precipitation Nowcasting

### v1

@@ Abstract
Precipitation nowcasting is an important spatio-temporal prediction task to predict the radar echoes sequences based on current observations, which can serve both meteorological science and smart city applications. Due to the chaotic evolution nature of the precipitation systems, it is a very challenging problem. Previous studies address the problem either from the perspectives of deterministic modeling or probabilistic modeling. However, their predictions suffer from the blurry, high-value echoes fading away and position inaccurate issues. The root reason of these issues is that the chaotic evolutionary precipitation systems are not appropriately modeled. Inspired by the nature of the systems, we propose to decompose and model them from the perspective of global deterministic motion and local stochastic variations with residual mechanism. A unified and flexible framework that can equip any type of spatio-temporal models is proposed based on residual diffusion, which effectively tackles the shortcomings of previous methods. Extensive experimental results on four publicly available radar datasets demonstrate the effectiveness and superiority of the proposed framework, compared to state-of-the-art techniques. Our code will be made publicly available soon.
@@ 1 Introduction
  • We develop a way to jointly train deterministic predictive model and probabilistic diffusion model in an end-to-end manner from the inspiration of precipitation chaotic evolution.
@@ 4 Overall Framework
Our framework, DiffCast, is designed to decompose the precipitation systems as the global motion trend and local stochastic residual, and utilizes a deterministic component and stochastic diffusion component to model them, respectively. The overview of our framework is illustrated in Figure 2 (a). Specifically, we apply a deterministic backbone as base predictor $\mathcal{P}_{\theta_{1}}$ to capture the global motion trajectory denoted as $\mu$. Note that any type of deterministic prediction models, e.g. SimVP [24], Earthformer [8], ConvGRU [27] etc., can be equipped into our framework without additional adaptation (Section 4.1). Then, the deterministic motion prior $\mu$ is leveraged for producing a residual sequence $r$ with ground truth $y$ to represent the local stochastics. We propose an auxiliary stochastic diffusion component to model the evolution of the residual distribution (Section 4.2). In the component, a sophisticated Global Temporal UNet(GTUnet) is designed to exploit the global motion prior, sequence segment consistency and inter-frame dependency for the evolution modelling (Section 4.3). Finally, the probabilistic predictive residual $\hat{r}$ and the deterministic prediction $\mu$ are added to form the final prediction $\hat{y}$.
@@ 4.2 Stochastic Residual Prediction
$p_{\theta_{2}}(r_{i}|\hat{r}_{i-1})=p(r^{T})\prod_{t=1}^{T}p_{\theta_{2}}(r_{i}^{t-1}|r_{i}^{t},\hat{r}_{i-1}),$ (9)
@@ 4.3 Global Temporal UNet (GTUNet)
Inter-frame temporal dependency. To better model the inter-frame dependency within a segment, our Global Temporal Unet is carefully designed with temporal attention blocks, as shown in orange part of Figure 3. It is indeed a variant of UNet in DDPM with temporal evolution. The temporal attention block, as shown in grey blue part of Figure 3, is constructed by concatenate operator, Resnet operator, normalization operator and temporal attention unit operator, followed by an up/downsample operator, which makes the predicted residual in the segment become temporal dependent. With the developed temporal evolution UNet structure, the inter-frame dependency can be effectively taken into account.
@@ 4.4 Training and Inference
For clarity, we summarize the computing flow of training and inference of our framework in Figure 2 (b). In the training phase, the deterministic predictor $\mathcal{P}_{\theta_{1}}$ first produces $\mu$. Then the residual $r$ is computed as Equation (7) and grouped into segments $s$ by Equation (17). Given a segment $s_{j-1}$, we predict the segment $s_{j}$ by our diffusion model $p_{\theta_{2}}$. Here the diffusion model is updated conventionally, namely sampling a step $t$, adding noise as scheduled, using the denoiser to predict noise and calculating loss as Equation (19) to update parameters $(\theta_{1},\theta_{2},\theta_{3})$. As for inference, the computing flow is similar with the only difference of diffusion part, where we first sample the $T$-state segment $s_{j}^{T}$ from $\mathcal{N}(0,I)$ and perform $T$-step denoising with $\epsilon_{\theta}$ to recover $\hat{s}_{j-1}$. Once all required residual segments are obtained we can compute the final prediction as Equation (13).
@@ 5.1 Experimental Setting
Data preprocess. As for all prediction sequences, we mainly focus on modeling the precipitation event. Hence, following [5], we separate the continuous sequence into multiple events for MeteoNet and Shanghai-Radar corpus. As mentioned in [38] that the increasing length of initial frames $L_{in}$ cannot provide substantial improvements for forecast, we set the challenging prediction task to predict 20 frames given 5 initial frames (i.e. $5\rightarrow 20$) except for the CIKM dataset, where only $5\rightarrow 10$ can be used due to its sequence length limitation. For all the datasets, we keep the original temporal resolution but downscale the spatial size to $128\times 128$, due to the limitation of our computation resource.
@@ 5.3 Analysis and Discussions
Table 2: Analysis of the necessarity of deterministic loss.
Method CSI CSI-pool4 CSI-pool16 HSS LPIPS SSIM
@@ 7 Datasets Detials
SEVIR, the Storm EVent ImagRy (SEVIR) [33] is an annotated, curated and spatio-temporally aligned dataset across five multiple data types including visible satellite imagery, infrared satellite imagery (mid-level water vapor and clean longwave window), NEXRAD radar mosaic of VIL(vertically integrated liquid mosaics) and ground lightning events. In this paper, we focus on the short term weather forecasting task and select all the radar mosaics of VIL as the main data. The dataset contains 20393 weather events from multiple sensors in 2017-2020. Each event consists of a 4-hour length sequence of images sampled in 5 minute steps covering 384 km×$\times$384 km patches sampled at locations throughout the continental U.S.. As our task is to predict the future VIL up to 20 frames (120 min) given 5 observed frames (30 min), we following [8] to sample the 25 continuous frames with stride = 12 in every event and split the dataset into training, validation and test sets with the time point January 1, 2019 and June 1, 2019, respectively. The frames are rescaled back to the range 0-255 and binarized at thresholds [16,74,133,160,181,219] to calculate the CSI and HSS following original settings in [33].
@@ 8 DiffCast: Implementation Details
In this section, we will give a detailed description of the implementation for DiffCast’s training and inference process and our experimental settings as well.
@@ 9 Additional Analysis
In this section, we give an extra analysis on the design of framework loss.
We decompose the determinism and local stochastics in precipitation evolution and model them with a deterministic component and a residual diffusion component, respectively. Different from other two-stage frameworks, we train the overall framework in an end-to-end manner to simulate the interplay between the determinism and uncertainty, which indicate that the gradient from stochastic diffusion denoising loss can also optimize the deterministic backbone. In Table 5, we compare the performance between the pure deterministic backbones and the intermediate output $\mu$ from deterministic component in DiffCast. The results show that the stochastic loss indeed leads to a positive optimization on most of the deterministic backbones.
Figure 8: CSI with threshold
@@
We point out that the conventional deterministic methods always under-estimate the high-value echoes with the increasing lead time (shown in Figure 1). To further investigate this, we show in Figure 8 the performance of DiffCast_SimVP, in terms of different thresholds. We observe that with the stochastic modeling the high-value echoes indeed can be more accurately maintained and predicted.

### v2

@@ Abstract
Precipitation nowcasting is an important spatio-temporal prediction task to predict the radar echoes sequences based on current observations, which can serve both meteorological science and smart city applications. Due to the chaotic evolution nature of the precipitation systems, it is a very challenging problem. Previous studies address the problem either from the perspectives of deterministic modeling or probabilistic modeling. However, their predictions suffer from the blurry, high-value echoes fading away and position inaccurate issues. The root reason of these issues is that the chaotic evolutionary precipitation systems are not appropriately modeled. Inspired by the nature of the systems, we propose to decompose and model them from the perspective of global deterministic motion and local stochastic variations with residual mechanism. A unified and flexible framework that can equip any type of spatio-temporal models is proposed based on residual diffusion, which effectively tackles the shortcomings of previous methods. Extensive experimental results on four publicly available radar datasets demonstrate the effectiveness and superiority of the proposed framework, compared to state-of-the-art techniques. Our code is publicly available at https://github.com/DeminYu98/DiffCast.
@@ 1 Introduction
  • Inspired by the natural meteorological mechanism, we propose to simultaneously train the deterministic predictive component and the stochastic diffusion component in an end-to-end manner and validate the effectiveness and necessity.
@@ 4 Overall Framework
Our framework, DiffCast, is designed to decompose the precipitation systems as the global motion trend and local stochastic residual, and utilizes a deterministic component and stochastic diffusion component to model them, respectively. The overview of our framework is illustrated in Figure 2 (a). Specifically, we apply a deterministic backbone as base predictor $\mathcal{P}_{\theta_{1}}$ to capture the global motion trajectory denoted as $\mu$. Note that any type of deterministic prediction models, e.g. SimVP [24], Earthformer [8], ConvGRU [27] etc., can be equipped into our framework without additional adaptation (Section 4.1). Then, the deterministic motion prior $\mu$ is leveraged for producing a residual sequence $r$ with ground truth $y$ to represent the local stochastics. We propose an auxiliary stochastic diffusion component to model the evolution of residual distribution (Section 4.2). In the component, a sophisticated Global Temporal UNet(GTUnet) is designed to exploit the global motion prior, sequence segment consistency and inter-frame dependency for the evolution modelling (Section 4.3). Finally, the probabilistic predictive residual $\hat{r}$ and the deterministic prediction $\mu$ are added to form the final prediction $\hat{y}$.
@@ 4.2 Stochastic Residual Prediction
$p_{\theta_{2}}(r_{i}^{0:T}|\hat{r}_{i-1})=p(r^{T})\prod_{t=1}^{T}p_{\theta_{2}}(r_{i}^{t-1}|r_{i}^{t},\hat{r}_{i-1}),$ (9)
@@ 4.3 Global Temporal UNet (GTUNet)
Inter-frame temporal dependency. To better model the inter-frame dependency within a segment, our Global Temporal UNet is carefully designed with temporal attention blocks, as shown in orange part of Figure 3. It is indeed a variant of UNet in DDPM with temporal evolution. The Temp-Attn Block, as shown in grey blue part of Figure 3, is constructed by concatenate operator, Resnet operator, normalization operator and temporal attention unit operator, followed by an up/downsample operator, which makes the predicted residual in the segment become temporal dependent. With the developed temporal evolution UNet structure, the inter-frame dependency can be effectively taken into account.
@@ 4.4 Training and Inference
For clarity, we summarize the computing flow of training and inference of our framework in Figure 2 (b). In the training phase, the deterministic predictor $\mathcal{P}_{\theta_{1}}$ first produces $\mu$. Then the residual $r$ is computed as Equation (7) and grouped into segments $s$ by Equation (17). Given a segment $s_{j-1}$, we predict the segment $s_{j}$ by our diffusion model $p_{\theta_{2}}$. Here the diffusion model is updated conventionally, namely sampling a step $t$, adding noise as scheduled, using the denoiser to predict noise and calculating loss as Equation (19) to update parameters $(\theta_{1},\theta_{2},\theta_{3})$. As for inference, the computing flow is similar with the only difference of diffusion part, where we first sample the $T$-state segment $s_{j}^{T}$ from $\mathcal{N}(0,I)$ and perform $T$-step denoising with $p_{\theta_{2}}$ to recover $\hat{s}_{j}$ given $\hat{s}_{j-1}$. Once all required residual segments are obtained we can compute the final prediction as Equation (13).
@@ 5.1 Experimental Setting
Data preprocess. As for all sequences, we mainly focus on modeling the precipitation event. Hence, following [5], we separate the continuous sequence into multiple events for MeteoNet and Shanghai-Radar corpus. As mentioned in [39] that the increasing length of initial frames $L_{in}$ cannot provide substantial improvements for forecast, we set the challenging prediction task to predict 20 frames given 5 initial frames (i.e. $5\rightarrow 20$) except for the CIKM dataset, where only $5\rightarrow 10$ can be used due to its sequence length limitation. For all the datasets, we keep the original temporal resolution but downscale the spatial size to $128\times 128$, due to the limitation of our computation resource.
@@ 5.2 Experimental Results
Table 2: Analysis of the necessarity of deterministic loss.
Method CSI CSI-pool4 CSI-pool16 HSS LPIPS SSIM
@@ 6 Conclusion
Acknowledgement
This work was supported in part by NSFC under Grants 62376072, 62272130, and in part by Shenzhen Science and Technology Program No. KCXFZ20211020163403005 and in part by Science and technology innovation team project of Guangdong Meteorological Bureau (GRMCTD202104), Innovation and Development Project of China Meteorological Administration (CXFZ2022J002), Shenzhen Hong Kong Macao science and technology plan project (SGDX20210823103537035).
@@ 7 Datasets Detials
SEVIR, the Storm EVent ImagRy (SEVIR) [33] is an annotated, curated and spatio-temporally aligned dataset across five multiple data types including visible satellite imagery, infrared satellite imagery (mid-level water vapor and clean longwave window), NEXRAD radar mosaic of VIL(vertically integrated liquid mosaics) and ground lightning events. In this paper, we focus on the short term weather forecasting task and select all the radar mosaics of VIL as the main data. The dataset contains 20393 weather events from multiple sensors in 2017-2020. Each event consists of a 4-hour length sequence of images sampled in 5 minute steps covering 384 km×$\times$384 km patches sampled at locations throughout the continental U.S.. As our task is to predict the future VIL up to 20 frames (100 min) given 5 observed frames (25 min), we follow [8] to sample the 25 continuous frames with stride = 12 in every event and split the dataset into training, validation and test sets with the time point January 1, 2019 and June 1, 2019, respectively. The frames are rescaled back to the range 0-255 and binarized at thresholds [16,74,133,160,181,219] to calculate the CSI and HSS following original settings in [33].
@@ 8 DiffCast: Implementation Details
In this section, we will give a detailed description of the implementation for DiffCast’s main architecture and its training and inference process, as well as our experimental settings.
Architecture of DiffCast. We have described the main architecture of the DiffCast model in section 4.3. Here, we present our detailed implementation of the Temp-Attn Block and GlobalNet, which are mainly composed of temporal attention operator[3, 30] and ConvGRU operator[27, 37], respectively, as summarized in Table 5.
Table 5: Detailed implementation of our Temp-Attn Block and GlobalNet.
Temp-Attn Block
ResBlock×$\times$2 2$\times$[Conv3x3 + GroupNorm8+ SiLU] + Conv3x3 Res Operator
Temporal Attention Conv5x5 (Spatial) + Conv1x1 (Temporal)+FC Attention Operator
Down/Upsampler Conv1x1
GlobalNet
ResBlock×$\times$4 2$\times$[Conv3x3 + GroupNorm8+ SiLU] + Conv3x3 Res Operator
ConvGRU×$\times$4 Conv3x3 + Conv3x3; HiddenState GRU Operator
Downsampler×$\times$4 Conv1x1
@@ 9 Additional Analysis
In this section, we give an extra analysis on the design of framework loss, the computational complexity and the hyperparameter $K$.
w/o stochastic Loss. We decompose the determinism and local stochastics in precipitation evolution and model them with a deterministic component and a residual diffusion component, respectively. Different from other two-stage frameworks, we train the overall framework in an end-to-end manner to simulate the interplay between the determinism and uncertainty, which indicate that the gradient from stochastic diffusion denoising loss can also optimize the deterministic backbone. In Table 6, we compare the performance between the pure deterministic backbones and the intermediate output $\mu$ from deterministic component in DiffCast. The results show that the stochastic loss indeed leads to a positive optimization on most of the deterministic backbones.
@@
Figure 8: CSI with threshold
Figure 9: Qualitative results of SimVP w/o stochastic loss.
We point out that the conventional deterministic methods always under-estimate the high-value echoes with the increasing lead time (shown in Figure 1). To further investigate this, we show in Figure 8 the performance of DiffCast_SimVP, in terms of different thresholds. We observe that with the stochastic modeling the high-value echoes indeed can be more accurately maintained and predicted. Additionally, we show the qualitative results of SimVP w/o stochastic diffusion loss in Figure 9. The results indicate that DiffCast__\__SimVP-$\mu$ alleviates the echo value fading away issue compared to SimVP, which implies that stochastic objectives indeed help optimize the deterministic model.
Complexity and Hyperparameter $K$. In Table 7, we report the tradeoff between model size, memory and time cost conditioned on different segment length based on our experimental setting with batchsize=4 for 30K training iterations. $K$ is selected from {$2,4,5,10$} on validation set and $K$=5 delivers the best tradeoff. There are more requirements for memory compared with base predictor but it is acceptable in our practical application. It is notable that the model size is not influenced by $K$.
Table 7: Complexity analysis and hyperparameter $K$.
Training Inference
Model Size CSI Memory Time Cost Memory Time Cost
MAU 20.13M 0.2463 21759MB 14.5h 2297MB 0.29s
DiffCast_MAU(K=2) 66.38M 0.2638 43815MB 22.8 h 3881MB 42s
DiffCast_MAU(K=4) 66.39M 0.2697 35471MB 20.5 h 3731MB 21s
DiffCast_MAU(K=5) 66.40M 0.2716 33791MB 19.5 h 3815MB 16s
DiffCast_MAU(K=10) 66.43M 0.2548 30443MB 18.5h 4065MB 8s